# Uni-MuMER: Unified Multi-Task Fine-Tuning of Vision-Language Model for Handwritten Mathematical Expression Recognition

### v2

@@ 7 Conclusion
This paper introduces Uni-MuMER, a unified multi-task fine-tuning framework for HMER that leverages the large-scale VLM Qwen-2.5-VL-3B. By fine-tuning it without any architectural changes, Uni-MuMER integrates Tree-aware Chain-of-Thought, Error-Driven Learning, and Symbol Counting tasks to jointly address structured spatial reasoning, error correction, and long-expression consistency in expression recognition. Experiments on the CROHME and HME100K datasets demonstrate that Uni-MuMER achieves SOTA accuracy, surpassing all prior approaches.
@@ Appendix A Prompt Design for Multi-Task Training
¯¯a (above)
¯    x (right)
¯¯        0 (sub)
¯    + (right)
¯    b (right)
¯    y (right)
¯¯        0 (sub)
¯    + (right)
¯    c (right)
¯    $\sqrt$ (below)
¯¯        a (above)
¯¯¯            2 (sup)
¯¯        + (right)
¯¯        b (right)
¯¯¯            2 (sup)
[…]
Example A.3:  Prompt for Error Detection Prompt: I have an image of a handwritten mathematical expression and its OCR recognition result. Please help me to detect possible errors in the recognition result and mark the places where errors occur with <error_start> <error_end> and <deleted>.
[…]
@@
[…]
Example A.4:  Prompt for Error Correction Prompt: I have an image of a handwritten mathematical expression and a predicted formula with error tags, correct the formula by modifying the parts marked with <error_start> and <error_end> and inserting content where <deleted> are present. Output the modifications in a single string using the following format: REPLACE:old -> new for errors to be replaced. INSERT:content after the position for missing content. DELETE:to_delete for parts to be removed.
[…]
@@
|: 2, $\frac$: 1, a: 2, x: 1, _: 2, 0: 2, +: 3, b: 2, y: 1, c: 1, $\sqrt$: 1, ^: 2, 2: 2
[…]
@@ B.1 Data Preprocess
[…]
@@
Total (with 4 tasks) 38112 38379 254335 739969 163983
@@
Task Data Construction  With the cleaned datasets in hand, we construct the multi-task training samples. Each image can yield multiple training samples, one for each task. The lower part of Tab. 6 lists the number of training samples per task for each dataset. For the Tree-CoT, Symbol Counting, and Vanilla HMER tasks, we typically generate one sample per image, so their counts per dataset are equal to the number of remaining images. For the Error Detection and Error Correction subtasks, we collect all error corpus generated by each dataset. Further inspired by recent work on incremental refinement LATTE [26], our error-correction task is specifically constrained to correcting one error per round, which is more effective than inputting the whole incorrect expression. In total, across all datasets, our training set comprises roughly 1.6 million samples derived from about 392k unique images, spanning the four tasks.
@@ B.2 Evaluation Datasets and Metrics
Evalution Metric  The Character Error Rate (CER)  CER is computed via Levenshtein distance at the Unicode character level, normalized by ground-truth length. It is widely used in online handwriting datasets (e.g., MathWriting) to detect localization errors, brace imbalances, and missing superscripts. EditScore  Following MathNet, we adopt EditScore, defined as 1Edit $1-\tfrac{\text{Edit distance}}{\max(|\hat{y}|,|y|)}$. EditScore, the sequence-level counterpart to CER, strongly correlates with human judgments in printed Mathematical Expression Recognition (MER) tasks and tolerates minor reordering and length variations while preserving semantics. BLEU-4  For consistency and comparability with existing corpus in the im2latexv2 domain, we incorporate the BLEU-4 metric, calculated based on 4-gram overlap after normalization, which effectively quantifies n-gram overlaps and remains a valuable metric.
@@ B.3 Evaluation Prompting Protocol
Uni-MuMER (Ours)   We used the Vanilla HMER prompt to elicit the final output when evaluating our fine-tuned model on the test sets. Instead of explicitly prompting the model to show its chain-of-thought or to count symbols during this evaluation phase, we directly request the final output. The rationale is that our multi-task training has already imbued the model with better internal reasoning and understanding; at test time, we just want the final answer, not the intermediate steps. Thus, for each test image, we simply provided the vanilla prompt like: "I have an image of a handwritten mathematical expression. Please write out the expression of the formula in the image using format." to the model, along with the handwritten expression image.
@@ B.4 Training Detail
Implementation Details   We fine-tune Qwen2.5-VL-3B using full-parameter update. Besides, we employ a cosine decay schedule to gradually reduce the learning rate to 10% of its initial value ($1\times 10^{-5}$) by the end of training. Training is conducted for a single epoch using all tasks and mixed dataset batches. Due to memory constraints, we adopt mixed-precision training, which enables us to fit a batch size of 512 within approximately 80 GB of VRAM per GPU (4×A800). This large batch size was specifically chosen to accommodate the diversity of tasks and stabilize the training process.
Training time  We trained for a single epoch over the combined multi-task dataset. One epoch in this context amounted to roughly 1.6 million training samples, since each image contributes multiple task-specific samples. Iterating over 1.6M samples with batch 512 yields about 3125 optimization steps. In practice, this single epoch took 78 hours to complete on the 4$\times$ A800 setup. Although this runtime slightly exceeds that of training lightweight specialized models from scratch, our approach achieves convergence to a high-performance level within just a single epoch.
@@ C.1 Effects of Data Diversity Compared with Specialized HMER Model
Table 7: Impact of incrementally training data scaling on ExpRate@CDM. Starting from the CROHME training set, we incrementally incorporate four additional datasets. The model is retrained from scratch and evaluated on CROHME-Average at each expansion step. Compared with CoMER, Vanilla (baseline), and Uni-MuMER.
@@ C.2 Inference Efficiency
We evaluated the inference speed and model size of our proposed Uni-MuMER model against the baseline methods CoMER and TAMER, utilizing a single NVIDIA A800 GPU. To ensure a consistent and fair comparison, we used the HME100K test dataset, comprising 24,607 real-scene handwritten mathematical expression images. Additionally, we leveraged vLLM, an optimized transformer-based inference engine known for accelerating decoding processes through efficient batch processing and caching mechanisms. The comparative efficiency metrics are summarized in Table 8. As demonstrated, Uni-MuMER significantly surpasses previous HMER models in terms of inference speed, achieving an FPS approximately eightfold higher than existing baselines under identical hardware conditions. As a result, Uni-MuMER is not only accurate but also efficient enough for practical use,
@@ D.1 Results on Handwritten Mathematical Expression
Tab. 9 is an extended version of Tab. 1. Specifically, Tab. 9 reports not only ExpRate but also the lenient metrics allowing up to $\leq 1$ and $\leq 2$ token errors. We compare Uni-MuMER against a broad range of baselines: open-source VLMs, closed-source VLMs, previous SOTA methods on HMER and their data augmentation versions.
Tab. 10 and Tab. 11 shows the results on the MNE dataset. This shows that the vanilla model is already competitive with the prior SOTA on these challenging sets. By introducing three data-driven tasks and external data, the Uni-MuMER and Uni-MuMER significantly outperform the previous best methods.
@@
Table 9: Performance comparison on CROHME datasets (%). We evaluate our model and four baselines using expression recognition rate (Exp), $\leq 1$, and $\leq 2$ on the CROHME 2014, 2016, and 2019 test sets. We use the following notation for reported results: All models are evaluated under zero-shot inference settings, a denotes training with data augmentation, p denotes reproduced by us, and denotes the use of external data.
@@
Exp\,$\uparrow$ $\leq 1\uparrow$ $\leq 2\uparrow$ Exp\,$\uparrow$ $\leq 1\uparrow$ $\leq 2\uparrow$ Exp\,$\uparrow$ $\leq 1\uparrow$ $\leq 2\uparrow$
@@
Exp\,$\uparrow$ @CDM \,$\uparrow$ CDM $\uparrow$ Exp\,$\uparrow$ @CDM\,$\uparrow$ CDM\,$\uparrow$ Exp\,$\uparrow$ @CDM\,$\uparrow$ CDM\,$\uparrow$
CoMER 59.73 - - 37.17 - - 24.04 - -
PosFormer 60.59 - - 38.82 - - 36.82 - -
MMHMER 62.03 - - 39.47 - - 34.29 - -
@@
Exp\,$\uparrow$ $\leq 1\uparrow$ $\leq 2\uparrow$ Exp\,$\uparrow$ $\leq 1\uparrow$ $\leq 2\uparrow$ Exp\,$\uparrow$ $\leq 1\uparrow$ $\leq 2\uparrow$
@@
Exp $\uparrow$ $\leq 1\uparrow$ $\leq 2\uparrow$CDM $\uparrow$ CDM $\uparrow$
@@
Uni-MuMER 71.93+1.78 87.08+1.27 92.24+0.99 73.67+1.87 97.33+0.36
Model Input CER \;$\downarrow$ Exp\;$\uparrow$ @CDM\;$\uparrow$ CDM\;$\uparrow$
PaLI Image 8.07 - - -
PaLI Ink 4.64 - - -
PaLI Ink + Image 4.55 - - -
PaLM-E Image 4.87 - - -
PaLM-E Ink 6.46 - - -
PaLM-E Ink + Image 4.22 - -
Uni-MuMER Image 4.00 51.45 69.11 94.84

### v3

@@ 7 Conclusion
This paper introduces Uni-MuMER, a unified multi-task fine-tuning framework for HMER , establishing a new paradigm by leveraging large-scale VLMs integrating multiple HMER tasks in a unified textual modality. Without architectural modifications, Uni-MuMER integrates Tree-aware Chain-of-Thought, Error-Driven Learning, and Symbol Counting tasks to jointly address structured spatial reasoning, error correction, and long-expression consistency in expression recognition, achieving super SOTA performance. This unified method provides enhanced accuracy, scalability, and faster inference speed, establishing a promising new research direction for future advancements in HMER.
8 Acknowledgement
This work is supported by the projects of Beijing Nova Interdisciplinary Program (20240484647) and National Natural Science Foundation of China (No. 62376012), which is also a research achievement of State Key Laboratory of Multimedia Information Processing and Key Laboratory of Science, Technology and Standard in Press Industry (Key Laboratory of Intelligent Press Media Technology).
@@ Appendix A Prompt Design for Multi-Task Training
ΨΨa (above)
Ψ    x (right)
$Ψ$        0 (sub)
Ψ    + (right)
Ψ    b (right)
Ψ    y (right)
$Ψ$        0 (sub)
Ψ    + (right)
Ψ    c (right)
Ψ    $\sqrt$ (below)
$Ψ$        a (above)
ΨΨΨ            2 (sup)
$Ψ$        + (right)
$Ψ$        b (right)
ΨΨΨ            2 (sup)
$\frac$ { a x _ { 0 } + b y _ { 0 } + c } { $\sqrt$ { a ^ { 2 } + b ^ { 2 } } } |
Example A.3: Prompt for Error Detection Prompt: I have an image of a handwritten mathematical expression and its OCR recognition result. Please help me to detect possible errors in the recognition result and mark the places where errors occur with ¡error_start¿ ¡error_end¿ and ¡deleted¿.
erroneous formula: — $\frac$ { a x _ { 0 } + b y _ { 0 } + c } { $\sqrt$ { d ˆ { 2 } + b ˆ { 2 } } } —
@@
— $\frac$ { a x _ { 0 } + b y _ { 0 } + c } { $\sqrt$ { ¡error_start¿ d ¡error_end¿ ˆ { 2 } + b ˆ { 2 } } } —
Example A.4: Prompt for Error Correction Prompt: I have an image of a handwritten mathematical expression and a predicted formula with error tags, correct the formula by modifying the parts marked with ¡error_start¿ and ¡error_end¿ and inserting content where ¡deleted¿ are present. Output the modifications in a single string using the following format: REPLACE:old -¿ new for errors to be replaced. INSERT:content after the position for missing content. DELETE:to_delete for parts to be removed.
Marked formula: — $\frac$ { a x _ { 0 } + b y _ { 0 } + c } { $\sqrt$ { ¡error_start¿ d ¡error_end¿ ˆ { 2 } + b ˆ { 2 } } } —
@@
—: 2, $\frac$: 1, a: 2, x: 1, _: 2, 0: 2, +: 3, b: 2, y: 1, c: 1, $\sqrt$: 1, ^: 2, 2: 2
— $\frac$ { a x _ { 0 } + b y _ { 0 } + c } { $\sqrt$ { a ˆ { 2 } + b ˆ { 2 } } } —
@@ B.1 Uni-MuMER-Data Preprocess
Data Cleaning  First, we exclude all instances containing CJK characters. For example, the HME100K dataset occasionally contains Chinese characters in annotations; these were removed to avoid confusing the model. Additionally, we remove characters or symbols that lack semantic value and negatively impact recognition accuracy, such as $\underline{\quad}$ and empty brace pairs ({}), which do not contribute to the math content.
@@
Total (with four tasks) 38112 38379 254335 739969 163983
@@
Task Data Construction  With the cleaned datasets in hand, we construct the multi-task training samples. Each image can yield multiple training samples, one for each task. The lower part of Tab. 9 lists the number of training samples per task for each dataset. For the Tree-CoT, Symbol Counting, and Vanilla HMER tasks, we typically generate one sample per image, so their counts per dataset are equal to the number of remaining images. For the Error Detection and Error Correction subtasks, we collect all error corpus generated by each dataset. Further inspired by recent work on incremental refinement LATTE [26], our error-correction task is specifically constrained to correcting one error per round, which is more effective than inputting the whole incorrect expression. In total, across all datasets, our training set comprises roughly 1.6 million samples derived from about 392k unique images, spanning the three tasks.
@@ B.2 Evaluation Datasets and Metrics
Evalution Metric  The Character Error Rate (CER)  CER is computed via Levenshtein distance at the Unicode character level, normalized by ground-truth length. It is widely used in online handwriting datasets (e.g., MathWriting) to detect localization errors, brace imbalances, and missing superscripts. EditScore  Following MathNet, we adopt EditScore, defined as $1-\tfrac{\text{Edit distance}}{\max(|\hat{y}|,|y|)}$. EditScore, the sequence-level counterpart to CER, strongly correlates with human judgments in printed Mathematical Expression Recognition (MER) tasks and tolerates minor reordering and length variations while preserving semantics. BLEU-4  For consistency and comparability with existing corpus in the im2latexv2 domain, we incorporate the BLEU-4 metric, calculated based on 4-gram overlap after normalization, which effectively quantifies n-gram overlaps and remains a valuable metric.
@@ B.3 Evaluation Prompting Protocol
Uni-MuMER (Ours)   We used the Vanilla HMER prompt to elicit the final output when evaluating our fine-tuned model on the test sets. Instead of explicitly prompting the model to show its chain-of-thought or to count symbols during this evaluation phase, we directly request the final output. The rationale is that our multi-task training has already imbued the model with better internal reasoning and understanding; at test time, we just want the final answer, not the intermediate steps. Thus, for each test image, we simply provided the vanilla prompt like: ”I have an image of a handwritten mathematical expression. Please write out the expression of the formula in the image using format.” to the model, along with the handwritten expression image.
@@ B.4 Training Detail
Implementation Details   We fine-tune Qwen2.5-VL-3B using a full-parameter update. Besides, we employ a cosine decay schedule to gradually reduce the learning rate to 10% of its initial value ($1\times 10^{-5}$) by the end of training. Training is conducted for a single epoch using all tasks and mixed dataset batches. Due to memory constraints, we adopt mixed-precision training, which enables us to fit a batch size of 512 within approximately 80 GB of VRAM per GPU (8×A800). This large batch size was specifically chosen to accommodate the diversity of tasks and stabilize the training process.
[…]
@@ C.1 Effects of Data Diversity Compared with Specialized HMER Model
Table 10: Effect of progressively adding training datasets on CROHME‑Avg ExpRate@CDM. At each step, models are trained from scratch on the cumulative data. Uni‑MuMER improves monotonically and consistently outperforms the Vanilla (single‑task) and CoMER baselines.
@@ C.2 Inference Efficiency
We evaluated the inference speed and model size of our proposed Uni-MuMER model against the baseline methods CoMER and TAMER, utilizing a single NVIDIA A800 GPU. To ensure a consistent and fair comparison, we used the HME100K test dataset, comprising 24,607 real-scene handwritten mathematical expression images. Additionally, we leveraged vLLM, an optimized transformer-based inference engine known for accelerating decoding processes through efficient batch processing and caching mechanisms. The comparative efficiency metrics are summarized in Table 11. As demonstrated, Uni-MuMER significantly surpasses previous HMER models in terms of inference speed, achieving an FPS approximately eightfold higher than existing baselines under identical hardware conditions. As a result, Uni-MuMER is not only accurate but also efficient enough for practical use.
C.3 Alternative explorations format
In developing the Tree-CoT and Symbol Counting tasks, we explored various output representation formats and their impact on model performance. Our goal was to find a format that is expressive enough to capture structure/counts but concise enough for the model to learn effectively. Here we report experiments with alternative formats for the expression tree serialization in Tree-CoT, as well as for the symbol counting task. We compare each alternative to the format ultimately used in Uni-MuMER (referred to as “our format”). All experiments in this section were conducted using only CROHME data for training (no external data), so that differences reflect format efficacy rather than additional data. We evaluate both in a Vanilla+X setting and in the full Uni-MuMER setting (all three tasks, but substituting the alternative format for the relevant task) to assess each format’s influence. Below, we outline each representation with an example.
[…]
The S-expression is merely the formula itself and adds no metadata. Its nested structure also hurts readability and organization. The JSON AST is excessively complex—it reflects JavaScript’s internal LaTeX format—and includes substantial redundancy. The “Simplified JSON AST” was manually trimmed and lacks a consistent standard. In contrast, our AST format (used in Uni-MuMER) strikes a balance between structural clarity and brevity with special delimiters to indicate hierarchical relationships. Our format avoids the heavy overhead of full JSON while preserving the necessary structure of the expression tree.
Tab. 12 reports the model performance using each of these tree formats, which shows that the choice of format significantly impacts accuracy. We attribute this gain to our format’s compactness and structured guidance: it provides clear structural cues (like JSON) but with far fewer extraneous tokens, making it easier for the decoder to learn the task.
Table 12: Performance impact of different expression tree formats for Tree-CoT. Bold numbers are best in column. Setup: Models are trained on the CROHME dataset and evaluated on CROHME 2014, 2016, and 2019 test sets and their average. Results are reported in terms of ExpRate and ExpRate@CDM. We compare the baseline Vanilla model against variants that incorporate the Tree-CoT task using various AST output formats, and likewise compare full Uni-MuMER models using those formats.
Model CROHME 2014 CROHME 2016 CROHME 2019 CROHME Average
Exp @CDM Exp @CDM Exp @CDM Exp @CDM
Vanilla 72.62 75.66 64.26 67.83 69.14 71.98 68.67 71.82
Vanilla + S-expression 72.21 75.56 66.87 70.97 68.06 71.73 69.05 72.75
Vanilla + Json AST 71.70 75.66 66.70 70.51 70.56 74.04 69.65 73.40
Vanilla + Simplified Json AST 72.62 76.47 66.96 70.18 69.89 73.06 69.82 73.24
Vanilla + our AST 73.33 75.76 68.00 70.97 71.23 74.40 70.85 73.71
Uni-MuMER (w/ S-expression) 74.85 78.07 68.07 73.14 72.56 76.14 71.83 75.78
Uni-MuMER (w/ Json AST) 75.46 79.51 70.10 74.28 72.98 76.06 72.85 76.62
Uni-MuMER (w/ Simplified Json AST) 74.44 78.27 70.36 74.63 73.90 77.48 72.90 76.79
Uni-MuMER (w/ our AST) 75.36 79.11 70.79 75.24 73.73 77.23 73.29 77.19
In addition, we investigated alternative formats for the Symbol Counting (SC) auxiliary task. The default format in Uni-MuMER expresses the symbol counts as a simple list of symbol-count pairs, separated by commas. We compared this with a more explicit map/dictionary format. For example, consider the expression x ˆ { 2 } + y ˆ { 2 } ¡ 1.
Example C.4: Map Counting format for Symbol Counting
{’x’: 1, ’^’: 2, ’2’: 2, ’+’: 1, ’y’: 1, ’<’: 1, ’1’: 1}
Example C.5: Our Counting format for Symbol Counting
x: 1, ^: 2, 2: 2, +: 1, y: 1, <: 1, 1: 1
The Map Counting format is semantically equivalent to ours but clearly more redundant (it introduces additional punctuation and quotation characters). We found that this extra overhead can make the sequence longer and potentially confuse the model (especially because symbols like braces and quotes have no direct meaning in the counting task itself). As a result, as shown in Tab. 13, the map format yielded slightly worse performance. The differences are modest, but consistent.
Table 13: Impact of Symbol Counting formats. Bold numbers are best in column. Setup: Models are trained on the CROHME dataset and evaluated on CROHME 2014, 2016, and 2019 test sets and their average. Results are reported in terms of ExpRate and ExpRate@CDM. We compare using a JSON-like Map Counting format versus our simple listing format for the SC task.
Model CROHME 2014 CROHME 2016 CROHME 2019 CROHME Average
Exp @CDM Exp @CDM Exp @CDM Exp @CDM
Vanilla 72.62 75.66 64.26 67.83 69.14 71.98 68.67 71.82
Vanilla + Map Counting 72.92 76.67 65.39 70.01 69.89 73.65 69.40 73.44
Vanilla + our Counting 71.91 76.85 67.04 72.28 70.64 75.15 69.86 74.76
Uni-MuMER (w/ Map Counting) 75.05 78.80 70.62 74.43 72.48 76.21 72.72 76.48
Uni-MuMER (w/ our Counting) 75.36 79.11 70.79 75.24 73.73 77.23 73.29 77.19
C.4 Hallucination for HMER
Large generative models are known to occasionally hallucinate output tokens that do not belong or omit tokens that should be present. In the context of HMER, both big VLMs and prior specialized models can produce spurious symbols. To quantitatively analyze this, we preliminarily adopt two metrics: Hallucination Token Rate (HTR) (percentage of predicted tokens not in the ground truth) and Missing Token Rate (MTR) (percentage of ground truth tokens not in the pred). Additionally, we categorized hallucinations by token type (numerical(0-9), variable(a-zA-Z), structural delimiters (e.g., [], {}, ()). This gives finer-grained error counts: HTR_num, HTR_var, HTR_delim for hallucinated numbers, letters, and delimiters, respectively. We evaluated these metrics across the CROHME 2014, 2016, and 2019 test sets, ensuring that no external data was used during training. The results are summarized below:
Table 14: Error analysis of hallucinations and omissions on CROHME. Results include ExpRate (Exp), ExpRate@CDM, overall HTR and MTR errors, as well as specific error types (HTR_num, HTR_var, HTR_delim). We compare two previous SOTA models ( CoMER and TAMER ), the Vanilla baseline, ablations adding each of our tasks individually (Tree-CoT, EDL, SC), and the full Uni-MuMER.
Model Exp @CDM HTR MTR HTR_num HTR_var HTR_delim
CoMER (SOTA’2022) 59.18 62.47 2.55 3.91 0.37 0.33 1.18
TAMER (SOTA’2024) 60.29 63.18 2.17 4.02 0.29 0.29 1.05
Vanilla (baseline) 68.64 71.74 4.08 2.42 1.08 0.29 1.68
Vanilla + Tree-CoT 70.85 73.71 1.67 1.56 0.14 0.21 0.85
Vanilla + EDL 70.30 73.76 2.58 1.94 1.09 0.22 0.83
Vanilla + CAN 69.86 74.76 2.15 2.06 0.23 0.27 1.08
Uni-MuMER (w/ 3 tasks) 73.29 77.19 1.18 1.39 0.12 0.15 0.60
Tab. 14 shows that our full model yields the highest accuracy while also having the lowest HTR and MTR of all methods. Besides, incorporating each auxiliary task individually already helps: adding the Tree-CoT task slashes hallucinations to 1.67% HTR, adding the Symbol Counting task alone significantly lowers omissions (MTR), and adding the Error-Driven Learning task reduces hallucinations of delim and var. These improvements suggest that Tree-CoT primarily mitigates structural hallucinations, EDL corrects symbol-level mistakes, and SC enforces global consistency.
C.5 CJK influence for HMER
Table 15: Examples of CJK‑containing image–caption pairs in HME100K that lie outside pure expression recognition, which influence the robustness of expression recognition and are outside the scope of math expression recognition.
image id caption
train_63827 = 88、5
train_11042 1、3、14 $\times$ (9.42 $\div$ 3.14)
train_70305 39的因数: 1, 3,
In Sec. 4.2, we excluded samples containing CJK characters of HME100K. This is primarily for two reasons:
  1. 1.
    Our prompt specifically targets recognizing mathematical expressions, not general textual content such as CJK characters. This filtering ensures a clear evaluation of expression recognition capability.
  2. 2.
    Some CJK-containing samples introduced problematic annotations unrelated to pure expressions (e.g., textual descriptions or numeric lists)
Having justified the filtering strategy, we additionally examine the effect of CJK characters. We first list typical CJK-containing captions to illustrate the distributional shift (Tab. 15), then compare performance under two evaluation settings: (i) test set filtered to remove CJK, and (ii) test set retaining CJK. We also test training variants that include or exclude CJK samples. For reference, we also include results for CoMER and TAMER (previous specialized models) evaluated on the same splits.
Table 16: Comparison of recognition performance on HME100K test sets with and without CJK characters. Results are reported in terms of ExpRate and ExpRate@CDM. We evaluate three models on each subset: CoMER (specialized, SOTA 2022), TAMER (specialized, SOTA 2024), and Uni-MuMER. For Uni-MuMER, we report two training conditions: w/o CJK and w/ CJK. For CoMER and TAMER, we evaluated open-sourced models under zero-shot inference settings.
Model HME100K Test w/ CJK HME100K Test w/o CJK
ExpRate ExpRate@CDM CDM ExpRate ExpRate@CDM CDM
CoMER 68.47 70.19 96.79 68.70 70.29 96.81
TAMER 69.50 71.12 97.02 69.74 71.25 97.03
Uni-MuMER (w/o CJK) 69.46 72.49 97.08 72.66 74.30 97.38
Uni-MuMER (w/ CJK) 71.63 73.60 97.32 70.98 74.32 97.38
As is shown in Tab. 16, the performance of CoMER and TAMER remained nearly identical before and after filtering. Our model still outperforms previous methods, with the CJK characters the same in the training set, which suggests minimal effect on comparability, and including vs. excluding the CJK samples in training had minimal effect on overall HMER accuracy.
We further examine whether including CJK characters in training has any noticeable effect on the cross-domain performance of Uni-MuMER. Tab. 16 and Tab. 17 summarizes ExpRate@CDM results for two model variants across benchmarks. These results indicate that adding the small set of CJK-containing samples introduces a tiny trade-off: it may slightly compromise performance on the main handwriting benchmarks, while providing a negligible benefit to certain out-of-domain cases. We conclude that excluding the CJK data is a reasonable choice to keep the training focused, with no meaningful loss in generality.
Table 17: Effect of including CJK during training evaluated on ExpRate@CDM across benchmarks. Results are reported on CROHME-Average, CROHME 2023 test, Mathwriting test, and Im2Latexv2 test. The best results for each column are highlighted in bold.
Model CROHME-Avg CROHME 2023 Mathwriting Im2Latexv2
Uni-MuMER (w/o CJK) 82.89 74.42 69.11 88.99
Uni-MuMER (w/ CJK) 82.21 72.14 68.91 90.28
@@ D.1 Results on Handwritten Mathematical Expression
Tab. 18 is an extended version of Tab. 1. Specifically, Tab. 18 reports not only ExpRate but also the lenient metrics allowing up to $\leq 1$ and $\leq 2$ token errors. We compare Uni-MuMER against a broad range of baselines: open-source VLMs, closed-source VLMs, previous SOTA methods on HMER, and their data augmentation versions.
Tab. 19 presents an extended evaluation of our method on the CROHME2023 test set, providing metrics including ExpRate, $\leq 1$ and $\leq 2$ token errors, as well as CDM-based evaluations. We benchmark Uni-MuMER and Uni-MuMER against various baselines. Our baseline model already demonstrates competitive performance against previous SOTA methods. Notably, after incorporating external data, Uni-MuMER exhibits a drop in ExpRate due to changes in LaTeX-style outputs but achieves outstanding performance in ExpRate@CDM.
Tab. 20 and Tab. 21 shows the results on the MNE dataset. This shows that the vanilla model is already competitive with the prior SOTA on these challenging sets. By introducing three data-driven tasks and external data, the Uni-MuMER and Uni-MuMER outperform the prior best methods.
@@
Table 18: Performance comparison on CROHME datasets (%). We evaluate our model and four baselines using expression recognition rate (Exp), $\leq 1$, and $\leq 2$ on the CROHME 2014, 2016, and 2019 test sets. We use the following notation for reported results: All results are zero-shot inference; a: data augmentation, p: reproduced results, : external data used.
@@
[…]
@@
Table 19: Performance on CROHME2023 Test set (%). Evaluated by ExpRate, ExpRate@CDM, and CDM. The denotes the use of external data. The p denotes the results reproduced by [55].
Method CROHME 2023
ExpRate $\uparrow$ $\leq 1\uparrow$ $\leq 2\uparrow$ ExpRate@CDM $\uparrow$ CDM $\uparrow$
VLPG 56.47 68.54 73.63      –      –
CoMER p 59.82 69.23 74.12      –      –
TST 60.72 70.51 75.00      –      –
TST 62.59 73.59 77.51      –      –
Vanilla (baseline) 68.39 79.13 87.57 69.24 94.43
Uni-MuMER 70.26+1.87 81.13+2.00 88.35+0.78 71.47+2.23 95.12+0.69
Uni-MuMER 70.00+1.61 79.78+0.65 87.00-0.57 74.42+5.18 95.33+0.90
@@
[…]
CoMER 59.73 37.17 24.04
PosFormer 60.59 38.82 36.82
MMHMER 62.03 39.47 34.29
@@
[…]
@@
[…]
PaLI Image 8.07
PaLI Ink 4.64
PaLI Ink + Image 4.55
PaLM-E Image 4.87
PaLM-E Ink 6.46
PaLM-E Ink + Image 4.22
Vanilla Image 4.15 50.51 66.46 94.20
Uni-MuMER Image 4.01 51.19 68.73 94.80
Uni-MuMER Image 4.00 51.45 69.11 94.84
@@
Exp $\uparrow$ $\leq 1\uparrow$ $\leq 2\uparrow$ @CDM $\uparrow$ CDM $\uparrow$
@@
Uni-MuMER 72.66+2.51 87.67+1.86 92.60+1.35 74.30+2.50 97.38+0.41